OpenStudy (anonymous):

A friend sent me this question. 2 times the square root of x minus the square root of x-5 = the square root of 2x-2 Solve for x

OpenStudy (jdoe0001):

\(\bf 2\sqrt{x}-\sqrt{x-5}=\sqrt{2x-2}\quad ?\)